OpenStudy (anonymous):

6x-3y=18 I forgot how to graph :(

alones (alones):

so let's solve this firsts for x. 6x−3y=18 Step 1: Add 3y to both sides. 6x−3y+3y=18+3y 6x=3y+18 Step 2: Divide both sides by 6. 6x6=3y+186 x=12y+3 here use this link put your equation there and it will show you how would it lok on the graph https://www.desmos.com/

OpenStudy (anonymous):

divide by 6 or multiply?

OpenStudy (anonymous):

you lost me on step 2

alones (alones):

OOH MY BAD;[ yea oka diviede 6x/6=3y+ WAIT LET ME DRAW THIS

alones (alones):

 THEN x=1/2y+3 HERE THAT LOOKS BETTE

OpenStudy (mathmale):

AloneS's approach is OK. You may find it easier to graph 6x-3y=18 by using the intercept method. First, let x=0. Then -3y=18 and y= ??? so one point to graph is (0, ? ) Next, let y=0. Then 6x=18. x= ? Then another point to graph is ( ? , 0). Draw a line thru the 2 points.

alones (alones):

Whew thanks mr @mathmale your steps are easier!!! :]

OpenStudy (anonymous):

so y=-6 and x=2?

OpenStudy (mathmale):

Before accepting these results, check them in the original equation. Does the point (0, -6) satisfy the equation? Does (2, 0) satisfy it? If so, your points are correct.

OpenStudy (anonymous):

how do i find the slope?

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

x=3 not 2 my bad

OpenStudy (mathmale):

Riley, I'm going to assume that you want the slope of the line you were drawing for this problem. Also, I'm assuming that x=3: (3,0) and that the other point is (0,-6). If these are wrong, change them, please. Steal info from these two points: let x0=3 and y0=0. Let x1=0 and y1=-6. Then the slope you want is found using this equation:\[\frac{ y1-y0 }{ x1-x0 }=m\]

OpenStudy (mathmale):

Try this calculation.